Maxey & Sons manufactures two types of storage cabinets—Type A and Type B—and applies manufacturing overhead to all units at the rate of $160 per machine hour. Production information follows. Type A Type B Anticipated volume (units) 32,000 60,000 Direct-material cost per unit $ 48 $ 72 Direct-labor cost per unit 53 53 The controller, who is studying the use of activity-based costing, has determined that the firm’s overhead can be identified with three activities: manufacturing setups, machine processing, and product shipping. Data on the number of setups, machine hours, and outgoing shipments, which are the activities’ three respective cost drivers, follow. Type A Type B Total Setups 180 140 320 Machine hours 64,000 90,000 154,000 Outgoing shipments 200 150 350 The firm’s total overhead of $24,640,000 is subdivided as follows: manufacturing setups, $5,376,000; machine processing, $14,784,000; and product shipping, $4,480,000.

Required: 1. Compute the unit manufacturing cost of Type A and Type B storage cabinets by using the company’s current overhead costing procedures.

2. Compute the unit manufacturing cost of Type A and Type B storage cabinets by using activity-based costing.

3. Is the cost of the Type A storage cabinet overstated or understated (i.e., distorted) by the use of machine hours to allocate total manufacturing overhead to production? By how much?

4. Assume that the current selling price of a Type A storage cabinet is $485.00 and the marketing manager is contemplating a $50 discount to stimulate volume. Is this discount advisable?
Answer #1

Requirement 1

Particulars Type A Type B
Direct Materials 48 72
Direct Labour 53 53
Manufacturing Overhead 320 240
Unit Product Cost 421 365

Supporting Work

Manufacturing Overhead for Type A = Machine Hours per Unit of Type A * Overhead Rate

Manufacturing Overhead for Type A = 2 * 160 = $ 320

Machine Hours per Unit for Type A = Total Machine Hours for Type A / Total Anticipated Units of Type A

Machu Hours per Unit for Type A = 64,000 / 32,000 = 2 Hours

Manufacturing Overhead for Type B = Machine Hours per Unit of Type B * Overhead Rate

Manufacturing Overhead for Type A = 1.5 * 160 = $ 240

Machine Hours per Unit for Type B = Total Machine Hours for Type B / Total Anticipated Units of Type B

Machu Hours per Unit for Type B = 90,000/ 60,000 = 1.5 Machine Hours

Requirement 2

As per Activity Based Costing

Particulars Type A Type B
Direct Materials 48 72
Direct Labour 53 53
Manufacturing Overhead 366.50 215.20
Unit Product Cost 467.50 340.20

Supporting Work

A B C = A/B
Particulars Overhead Costs Cost Driver Activity Allocation Rates
Manufacturing Setups 53,76,000 320 16,800 per Set-up
Machine Processing 1,47,84,000 154,000 96 per Machine Hour
Product Shipping 44,80,000 350 12,800 per Shipment
Total Overhead 2,46,40,000

Type A

Particulars Activity Allocation Rates Activity Level Total Overhead Costs
Manufacturing Setps 16,800 180 30,24,000
Machine Processing 96 64,000 61,44,000
Product Shipping 12,800 200 25,60,000
Total Overhead Costs 1,17,28,000

Cost per Unit = 1,17,28,000 / 32,000 = $ 366.50

Type B

Particulars Activity Allocation Rates Activity Level Total Overhead Costs
Manufacturing Setps 16,800 140 23,52,000
Machine Processing 96 90,000 86,40,000
Product Shipping 12,800 150 19,20,000
Total Overhead Cost 1,29,12,000

Manufacturing Overhead per Unit of B = 129,12,000/ 60,000 = $ 215.20

Formula for Total Overhead Cost = Activity Allocation Rates * Activity Level

Requirement 3

For Type A

Particulars Amount
Costs as per Traditional Method 421
Costs as per Activity Based Costing 467.50
Overcasted (Undercasted) (46.50)

As per Machine Hours Allocation Method Manufacturing Cost per Unit of Type A is Undercasted by $ 46.5 per Unit

For 32,000 Units = 32,000 * 46.5 = $ 14,88,000

Requirement 4

No the discount is not advisable because as per Activity Based Costing the per Unit of ech Unit of Type A is $ 467.50 whereas the discount will make the per unit Selling Price to $ 435 per Unit which is well below the actual cost per unit which in turn will provide the losses for the company so the manager should not make a decision to provide discount on the Product.